17. A farmer has 100 acres of tillable land on which corn, tomatoes, beans, peas, and carrots can be planted. The labor requirements, plant costs, and gross income for I acre of each of these crops are as follows: Labor Gross (worker-hours) Costs (S) Income (S) Corn 95 1300 275 345 435 20 200 Tomatoes Beans Peas Carrots 120 25 35 40 75 nd The farmer has available up to 3600 hr of labor at $7.25/hr. However, the labor and plant costs must be paid before any income from the sale of the produce is realized. The farmer has $3000 in capital to invest in this years planting and can borrow up to another $12,000 if desired. Any such loan would have a 9% annual interest rate but would be repaid within 4 months. Land unused for any of these vegetables must be maintained by planting ground cover. One acre of cover crops requires 2 hr of labor and costs $9. Determine a planting schedule that maximizes net income
Answer #1
Labor charges $7.25/Hr
Capital available $3,000
Available Labor Hrs
Cover crops cost $9 / 2 Hr
Total cost on labour with $3000 = 3000/7.25 = 400
Crop: Total labor Hrs: Costs ($): Gross Income ($): Net Income per Acre ($): % age income: Net Income per Labor Hr: Wt. Avg. Net Income per Acre Per labor Hr: Resource Allocation (of 400 available Labor Hrs): Net Income at 400 Labor Hrs allocation ($): Resource Allocation (of 1705 borrowed Lab Hrs: Net Income at 1705 Labor Hrs ($12,360) allocation ($): Resource Allocation of 3600 available Hrs: Net Income at 3600 available Labor Hrs ($):
Corn 5 20 95 75 3.64 15.00 29.60 118.4 1,776.00 504.68 7,570.20 1065.60 15984
Tomatoes 120 200 1300 1100 53.40 9.17 18.09 72.36 663.30 308.43 2,827.32 651.24 5969.7
Beans 25 55 275 220 10.68 8.80 17.36 69.44 611.07 295.99 2,604.69 624.96 5499.648
Peas 35 40 345 305 14.81 8.71 17.19 68.76 599.19 293.09 2,554.07 618.84 5392.74857
Carrots 40 75 435 360 17.48 9.00 17.76 71.04 639.36 302.81 2,725.27 639.36 5754.24
2060 50.68 4,288.93 18,281.55 3600.00    182,451.43
Borrowing $12,000 at 9% annualized Interest for 4 months (4/12) will cost, overall = $12,000 x 0.09/3 = $12,360
$12,360 can bring 12,360/7.25 Labor Hrs = 1705 Hrs
Net Income on Available capital Investment = $4,288.93 - $3,000 = $1,288.93
Net Income on Borrowed capital Investment = $18,281.55 - $12,360 = $1,288.93 = $5,921.55